# Costless correction of chain based nested sampling parameter estimation in gravitational wave data and beyond

### v1

@@ 3.3 Likelihood binning method
In this first approach, we bin the phantom points from the run by their likelihood values, such that each phantom point is assigned to the dead point to which it is closest in log-likelihood. In this way, each dead point is now associated with a set of points which sit very close to the contour defined by it; we make the assumption that, though the phantom points do not lie exactly on the dead point’s iso-likelihood contour, they are still representative of the distribution of the $f(\theta)$ values along the contour. Then, for each dead point in our sum in equation 8, we resample a new $f(\theta)$ value from the associated bin, which includes the original dead point itself, as well as sampling a new $\Delta X$. The error estimate is then obtained by repeating this process many times and taking the standard deviation of the resulting distribution of estimates.
Figure 4: For each of the first 10 nested sampling runs performed on our simulated dataset, the new distribution of chirp mass estimates are plotted, obtained from resampling both shrinkage ratios and chirp mass values from likelihood bins around each contour (top). Resampling both the weights and the chirp mass gives wider error bars, as expected, and even by eye these seem more consistent with the spread of estimates across runs. Testing the likelihood binning method more rigorously, we see that the percentiles of the overall mean chirp mass are now consistent with being uniformly distributed (bottom, green). This indicates that the true variation of the chirp mass along the contours is now being correctly accounted for, and thus the stochasticity of nested sampling parameter estimation is properly captured.
@@
In reality, the phantom points are slightly correlated, which can lead to biased results due to the points not independently populating the prior. However, the large majority of the correlation can be removed by only using every, for example, $5^{\textrm{th}}$ point in the chains, as the correlation in the chirp mass becomes negligible after a few steps. The results for the chirp mass are shown in Figure 4. As demonstrated by the K-S test $p$-value and the histogram, the distribution of percentiles is now much more consistent with a uniform distribution. This shows that the error bars now accurately capture the variation in the chirp mass over the contours, and reflect the run-to-run differences from nested sampling due to this additional stochasticity.
3.4 Reconstructed runs method
@@
In order to explore the second method, we first note that there is nothing wrong with any of the phantom points in terms of their suitability for use in a nested sampling run, except that we have already chosen to use another point (the associated dead point) which may be too correlated with the phantom point to use both. Hence, we may take the $1^{\textrm{st}}$ phantom point in every chain in the run and combine these carefully to form an equally valid nested sampling run to the original. Thus, from every run we are able to reconstruct multiple equally valid ‘phantom runs’. It is true that some of these runs will be correlated to each other, but, as before, this correlation can be largely removed by only using a subset of the phantom points.
@@
These extra phantom runs are akin to performing multiple nested sampling runs on the same dataset, but, crucially for gravitational wave inference, at no extra computational cost. The parameter estimates from each of these reconstructed runs, as well as the original run, can then be combined and the corresponding new error bar may be computed from this. The resulting percentiles plot is shown in Figure 5, where it can be seen that, as with the first method, the new error bars are now much more consistent with the spread of estimates across multiple runs.
3.5 Verifying accuracy of error bar
For certain parameters, the chain length of the sampler may not be long enough to accumulate enough uncorrelated phantom points that populate the prior in the correct way. In this case, the new error bars from the above methods, though certainly wider than those from the usual method, may still not be wide enough to capture the full variation of the parameter along the contours. For gravitational waves, parameters like the mass ratio and the luminosity distance may suffer from this. If one is able to perform many nested sampling runs on the dataset, the error bars can be checked rigorously in a manner similar to the percentiles analysis above. However, this is inefficient and in many cases simply not viable. We need to be able to know from a single run whether the new error bars are now fully correct for a given parameter. Below, we present a method to check for convergent results of the parameter estimates, indicating whether the error bars have been quantified correctly.
For the default chain length in PolyChord the phantom points are very well distributed in their chirp mass values in order to be able to obtain an accurate error bar for this parameter. But they are more correlated in their luminosity distance values, meaning that the error bars obtained for this parameter will still be an underestimate of the true error. We need to employ a longer chain length in the run in order to obtain sufficient usable phantom points for the two methods described above. We can check whether there are enough uncorrelated phantom points for a given parameter by dividing the points into two halves, according to their position within the chains. We can then take each half separately and use either the likelihood binning method or the reconstructed runs method to obtain a set of parameter estimates, checking whether the two sets of estimates are in agreement using the Kolmogorov-Smirnov 2-sample test.
@@ 3.5 Verifying accuracy of error bar
Figure 6: The chirp mass was estimated from a single run using the likelihood binning method described in Section 3.3. Two sets of estimates were produced, one using only the first ten phantom points in all the chains and the other using only the last ten. The resulting distributions are in agreement with each other, with the K-S 2-sample test statistic being $0.14$, with a $p$-value of $0.28$. This indicates that the phantom points were well distributed in the chirp mass and give an accurate estimate of the true error bar.
@@
Figure 7: The same analysis was repeated for the luminosity distance, a parameter on which our methods do not perform well with the default PolyChord chain length. This time the K-S 2-sample test statistic was $0.82$, with a $p$-value of $4.11\times 10^{-34}$. It is likely that the phantom points are still a little too correlated to their associated dead point in luminosity distance, and a longer chain length is needed as more of the points in the chain must be discarded.
@@
This is demonstrated in Figures 6 and 7. For the chirp mass, it can be seen that the estimates obtained from using only the first ten phantom points in the chain agree well with the estimates from the last ten phantom points in the chain. The same cannot be said for the luminosity distance, indicating that we may need a chain length longer than 20 in order to obtain enough uncorrelated phantom points to estimate the error on this parameter. It is important to note here that this does not mean the chain length of the run wasn’t long enough to produce uncorrelated posterior samples in the luminosity distance, only that it was not long enough to produce sufficient suitable phantom points.
@@
To examine this further, we performed a run with a chain length of $100$, to see how many phantom points must be discarded before we have enough uncorrelated points to provide convergent estimates of the luminosity distance. First, the estimates obtained from the first 20 phantom points in the chain are compared to the those from the last 80 points, shown in Figure 8. The two sets of estimates are very different and this is a clear indication that the original chain length of $20$ was insufficient to accurately estimate the luminosity distance error bars. Next, the estimates obtained from phantoms points at positions in the chains between 50 and 75 were compared to estimates from the last 25 points in the chains. The two sets of estimates in Figure 9 are now in agreement and show that the last 50 phantom points in the chain may be used in the likelihood binning or reconstructed runs method to provide an accurate estimate of the error bars.
@@
Figure 8: For a run with a chain length of $100$, the luminosity distance estimates from the first $20$ points in the chain are compared to those from the rest of the phantom points. The K-S 2-sample test statistic for the resulting set of estimates is $0.79$, with an accompanying $p$-value of $3.05\times 10^{-31}$. This indicates that the first 20 phantom points are not well distributed enough to make an accurate estimate of the luminosity distance error bars and more points are needed.
@@
Figure 9: The same analysis is repeated, but comparing estimates obtained from the phantom points in the chains between positions 50 and 75 and the estimates from the just the last 25 points in each chain. These two sets of estimates now show good agreement, with a test statistic of $0.19$ and a $p$-value of $0.05$. Using the last 50 phantom points in the chains for either method described in this paper would therefore give an accurate estimate of the true luminosity distance error bar.
4 Comparison to Higson method
@@
Figure 10: The Higson method (purple) and the methods presented in this paper (green) perform similarly well on the simulated BBH signals, producing percentile plots that are consistent with a uniform distribution. Both are a significant improvement on the usual ‘simulated weights’ method, and are distinct in their approaches so may be used in conjunction with each other.
@@
In order to compare the methods presented above to those obtained using the bootstraps method presented in Higson et al. (2018a), we use the nestcheck (Higson et al., 2018c) package to apply the Higson method on the same set of simulated BBH signals. In this method, a nested sampling run is split up into single live point runs, and then $n_{\textrm{live}}$ runs are sampled with replacement from these. This is repeated many times to build up a distribution of parameter estimates from which the error bar is estimated.
@@
Figure 10 shows the resulting percentiles plot for the chirp mass; the Higson method performs similarly to the phantom methods and provides an independent way to correctly estimate the error bars. On parameters where the phantom methods do not work as well for the default settings, such as the luminosity distance, the Higson method also does not produce wide enough error bars. The advantage of using phantom points is that they can confirm from a single run whether or not the error bars are correct for a given parameter, as described in the previous section.
@@
5 Coverage and Credible Intervals
@@
The additional error in nested sampling parameter estimation described in Section 3 affects not only our estimates of parameter means, but also the credible intervals. If a parameter inference procedure is accurate, we expect that the percentiles of the true parameters across many analysed datasets should be uniformly distributed (Talts et al., 2020; Samantha R Cook & Rubin, 2006). The tool typically used to check this is a $p$-$p$ plot (Morisaki & Raymond, 2020; Veitch et al., 2015b; Pankow et al., 2015; Smith et al., 2020; Berry et al., 2015; Del Pozzo et al., 2018; Romero-Shaw et al., 2020; Biwer et al., 2019; Sidery et al., 2014; Green & Gair, 2020), constructed from $N$ runs on injected signals where the true parameters are known.
@@
A $p$-$p$ plot shows the fraction of events for which the true parameter values lie within a given credible interval as a function of the credible interval. In the ideal case where the parameter inference procedure leads to the correct coverage, this plot should be a diagonal line. There is a statistical error that arises from the finite number, $N$, of datasets analysed, often indicated by a gray region around the ideal diagonal line (as in Figures 11 and 12), and this is usually the dominant source of error in $p$-$p$ plot analyses. However, as $N$ increases, the nested sampling error becomes important. This error can be evaluated using either the likelihood binning or reconstructed runs method, and we can place a ‘nested sampling confidence interval’ on our $p$-$p$ plot to account for the variation of the calculated percentiles from run to run on the same set of events.
@@
Figures 11 and 12 were produced using the likelihood binning method. For each posterior sample from a given event, the individual parameter values were resampled from bins of neighbouring phantom points. These resampled parameter values were then used to calculate a set of credible intervals for that event, from which the error on the credible interval was estimated. In Figure 11, a total of $200$ injected signals were analysed, with the injection parameters drawn from the prior, and the gray region shows the 3-$\sigma$ ($99.7$%) confidence interval for $N=200$ due to a finite event sample size. The purple region shows the corresponding 3-$\sigma$ confidence interval on the $p$-$p$ curve due to the nested sampling parameter estimation error. In scenarios where the $p$-$p$ plot lies outside of the gray region, using the nested sampling error bars may help to assess how much of the deviation is simply due to the stochastic nature of nested sampling. Figure 12 is produced from $1000$ injected signals, where now the statistical error for $N=1000$ and the nested sampling parameter estimation error are of similar sizes. Here, it is especially useful to consider the latter error in assessing to what extent the $p$-$p$ plot displays the expected coverage. Ideally, the gray confidence intervals should include both sources of error.
@@
Figure 11: The $p$-$p$ plot obtained for the chirp mass using bilby and PolyChord is shown. As with the parameter mean estimates, we only sample over four parameters, simply setting the rest to the injected value. The plot shows the cumulative distribution function (CDF) of the percentiles of the true chirp mass for $200$ events, calculated in the usual way, with the associated 3-$\sigma$ confidence interval due to the statistical error from the finite number of events (gray). The purple region shows the 3-$\sigma$ confidence interval on the calculated CDF of the percentiles due to the stochastic nature of nested sampling. The binomial error still dominates, but the nested sampling parameter estimation error can be useful in assessing the source of deviations from the expected distribution.
@@
Figure 12: The CDF of the percentiles of the true chirp mass is shown for $1000$ events, along with the 3-$\sigma$ confidence interval due to a finite event sample size (gray). The 3-$\sigma$ confidence interval due to the nested sampling parameter estimation error is also shown (purple). These are now of a comparable size, and in order to make an accurate assessment of the coverage, it is necessary to account for both sources of error.

### v2

@@ 3.3.1 Method 1: Likelihood binning
In this approach, we bin the phantom points from the run by their likelihood values, such that each phantom point is assigned to the dead point to which it is closest in log-likelihood. Thus, each dead point is now associated with a set of points which sit very close to the contour defined by it; we make the assumption that, though the phantom points do not lie exactly on the dead point’s iso-likelihood contour, they are still representative of the distribution of the $f(\theta)$ values along the contour. Then, for each dead point in our sum in equation 8, we resample a new $f(\theta)$ value from the associated bin, which includes the original dead point itself, as well as sampling a new $\Delta X$. The uncertainty estimate is obtained by repeating this process many times and taking the standard deviation of the resulting distribution of estimates.
@@
In reality, the phantom points are slightly correlated, which can lead to biased results due to the points not independently populating the prior. However, the large majority of the correlation can be removed by only using every, for example, $5^{\textrm{th}}$ point in the chains, as the correlation in the chirp mass becomes negligible after a few steps. This is discussed further in Section 4.1.
3.3.2 Method 2: Reconstructed runs
@@
In the second method, we first note that there is nothing wrong with any of the phantom points in terms of their suitability for use in a nested sampling run, except that we have already chosen to use another point (the associated dead point) which may be too correlated with the phantom point to use both. Hence, we may take the $1^{\textrm{st}}$ phantom point in every chain in the run and combine these carefully to form an equally valid nested sampling run to the original. Thus, from every run we are able to reconstruct multiple equally valid ‘phantom runs’. It is true that some of these runs will be correlated to each other, but, as discussed above, this correlation can be largely removed by only using a subset of the phantom points.
These extra phantom runs are akin to performing multiple nested sampling runs on the same dataset, but, crucially for gravitational wave inference, at no extra computational cost. The parameter estimates from each of these reconstructed runs, as well as the original run, can then be combined and the corresponding new error bar may be computed from this.
3.3.3 Verifying accuracy of error bar
For certain parameters, the chain length of the sampler may not be long enough to accumulate enough uncorrelated phantom points that populate the prior in the correct way. In this case, the new error bars from the above two methods, though certainly wider than those from the usual method, may still not be wide enough to capture the full variation of the parameter along the contours. For gravitational waves, parameters like the mass ratio and the luminosity distance may suffer from this. If one is able to perform many nested sampling runs on the dataset, the error bars can be checked rigorously, as in Section 3.1. However, this is inefficient and in many cases simply not viable. We need to be able to verify from a single run whether the new error bars are now fully correct for a given parameter. We propose a method to check for convergent results of the parameter estimates, indicating whether the error bars have been quantified correctly.
For the default chain length in PolyChord the phantom points are very well distributed in their chirp mass values in order to be able to obtain an accurate error bar for this parameter using them. But they are more correlated in their luminosity distance values, meaning that the error bars obtained for this parameter will still be an underestimate of the true uncertainty. We need to employ a longer chain length in the run in order to obtain sufficient usable phantom points for the two methods described above. We can check whether there are enough uncorrelated phantom points for a given parameter by dividing the points into two halves, according to their position within the chains. We can then take each half separately and use either the likelihood binning method or the reconstructed runs method to obtain a set of parameter estimates, checking whether the two sets of estimates are in agreement using the Kolmogorov-Smirnov 2-sample test. We demonstrate this method on the chirp mass and the luminosity distance in Section 4.3.
4 Results and Discussion
4.1 Likelihood binning
To demonstrate the efficacy of the likelihood binning method, we take the same 100 nested sampling runs used in Section 3.1, and apply the method as laid out in Section 3.3.1 to estimate the mean chirp mass of the simulated BBH event and its associated uncertainty. The results are shown in Figure 4. As in Figure 3, we include a histogram of the percentiles of the ‘true chirp mass’, estimated from the mean chirp mass of the 100 runs. As demonstrated by the K-S test $p$-value, the distribution of percentiles is now much more consistent with a uniform distribution. This shows that the error bars now accurately capture the variation in the chirp mass over the contours, and reflect the run-to-run differences from nested sampling due to this additional stochasticity.
Figure 4: For each of the first 10 nested sampling runs performed on our simulated dataset, the new distribution of chirp mass estimates are plotted, obtained from resampling both shrinkage ratios and chirp mass values from likelihood bins around each contour (top). Resampling both the weights and the chirp mass gives wider error bars, as expected, and even by eye these seem more consistent with the spread of estimates across runs. Testing the likelihood binning method more rigorously, we see that the percentiles of the overall mean chirp mass are now consistent with being uniformly distributed (bottom, purple). This indicates that the true variation of the chirp mass along the contours is now being correctly accounted for, and thus the stochasticity of nested sampling parameter estimation is properly captured.
In Section 3.3.1, we suggested using only every $5^{\textrm{th}}$ phantom point in order to mitigate the slight correlation between successive phantom points in a chain. In general, how many phantom points should be skipped will depend on the exact sampler and settings, but there will likely be several sensible choices that give comparable results. The metrics provided in this paper, particularly in Sections 3.3.3 and 4.3, provide a way to decide on a suitable choice. For the PolyChord sampler and the settings we chose for these runs, we found every $5^{\textrm{th}}$ point to be a sensible choice.
4.2 Reconstructed runs
As above, we use the same set of 100 nested sampling runs, this time applying the method described in Section 3.3.2. The resulting percentiles plot is shown in Figure 5, where it can be seen that, as with the first method, the new error bars are now much more consistent with the spread of estimates across multiple runs. The efficacy of these methods were also tested on other parameters besides the chirp mass, including parameters which were not explicitly sampled over. These results are included in Appendix A.
@@ 4.2 Reconstructed runs
4.3 Accuracy of new error bars
@@
Figure 6: The chirp mass was estimated from a single run using the likelihood binning method described in Section 3.3.1. Two sets of estimates were produced, one using only the first ten phantom points in all the chains and the other using only the last ten. The resulting distributions are in agreement with each other, with the K-S 2-sample test statistic being $0.13$, with a $p$-value of $0.37$. This indicates that the phantom points were well distributed in the chirp mass and give an accurate estimate of the true error bar.
@@
Figure 7: The same analysis was repeated for the luminosity distance, a parameter on which our methods do not perform well with the default PolyChord chain length. This time the K-S 2-sample test statistic was $0.77$, with a $p$-value of $1.92\times 10^{-29}$. It is likely that the phantom points are still a little too correlated to their associated dead point in luminosity distance, and a longer chain length is needed as more of the points in the chain must be discarded.
We illustrate below the method proposed in Section 4.3 to check whether the error bars from the likelihood binning method and the reconstructed runs method are indeed correct, without the need to perform multiple runs and plot a histogram of percentiles as above. From this section onwards, for conciseness all plots use the likelihood binning method only, but both methods perform similarly well on all considered examples.
@@
The two uncertainty estimation methods demonstrated above perform very well on parameters such as the chirp mass, but underestimate the uncertainty on parameters such as the luminosity distance. We apply the approach proposed in Section 4.3 to verify the error bars on both these parameters from a single run and obtain Figures 6 and 7. For the chirp mass, it can be seen that the estimates obtained from using only the first ten phantom points in the chain agree well with the estimates from the last ten phantom points in the chain. The same cannot be said for the luminosity distance, indicating that we may need a chain length longer than 20 in order to obtain enough uncorrelated phantom points to estimate the uncertainty on this parameter. It is important to note here that this does not mean the chain length of the run wasn’t long enough to produce uncorrelated posterior samples in the luminosity distance, only that it was not long enough to produce sufficient suitable phantom points.
@@
To examine this further, we performed a run with a chain length of $100$, to see how many phantom points must be discarded before we have enough uncorrelated points to provide convergent estimates of the luminosity distance. First, the estimates obtained from the first 20 phantom points in the chain are compared to those from the last 80 points, shown in Figure 8. The two sets of estimates are very different and this is a clear indication that the original chain length of $20$ was insufficient to accurately estimate the luminosity distance error bars. Next, the estimates obtained from phantoms points at positions in the chains between 50 and 75 were compared to estimates from the last 25 points in the chains. The two sets of estimates in Figure 9 are now in agreement and show that the last 50 phantom points in the chain may be used in the likelihood binning or reconstructed runs method to provide an accurate estimate of the error bars.
@@
Figure 8: For a run with a chain length of $100$, the luminosity distance estimates from the first 2020 points in the chain are compared to those from the rest of the phantom points. The K-S 2-sample test statistic for the resulting set of estimates is $0.81$, with an accompanying $p$-value of $3.94\times 10^{-33}$. This indicates that the first 20 phantom points are not well distributed enough to make an accurate estimate of the luminosity distance error bars and more points are needed.
@@
Figure 9: The same analysis is repeated, but comparing estimates obtained from the phantom points in the chains between positions 50 and 75 and the estimates from the just the last 25 points in each chain. These two sets of estimates now show good agreement, with a test statistic of $0.17$ and a $p$-value of $0.11$. Using the last 50 phantom points in the chains for either method described in this paper would therefore give an accurate estimate of the true luminosity distance error bar.
4.4 Comparison to Higson method
@@
Figure 10: The Higson method (green) and the methods presented in this paper (purple) perform similarly well on the simulated BBH signals, producing percentile plots that are consistent with a uniform distribution. Both are a significant improvement on the usual ‘simulated weights’ method, and are distinct in their approaches so may be used in conjunction with each other.
@@
In order to compare the methods presented in this paper to those obtained using the bootstraps method presented in Higson et al. (2018a), we use the nestcheck (Higson et al., 2018c) package to apply the Higson method on the same set of simulated BBH signals. In this method, a nested sampling run is split up into single live point runs, and then $n_{\textrm{live}}$ runs are sampled with replacement from these. This is repeated many times to build up a distribution of parameter estimates from which the error bar is estimated.
@@
Figure 10 shows the resulting percentiles plot for the chirp mass; the Higson method performs similarly to the likelihood binning method and provides an independent way to correctly estimate the error bars. On parameters where the phantom methods do not work as well for the default settings, such as the luminosity distance, the Higson method also does not produce wide enough error bars. The advantage of using phantom points is that they can confirm from a single run whether or not the error bars are correct for a given parameter, as demonstrated in the previous section.
@@
4.5 Coverage and credible intervals
@@
The additional uncertainty in nested sampling parameter estimation described here affects not only our estimates of parameter means, but also the credible intervals. If a parameter inference procedure is accurate, we expect that the percentiles of the true parameters across many analysed datasets should be uniformly distributed (Talts et al., 2020; Samantha R Cook & Rubin, 2006). The tool typically used to check this is a $p$-$p$ plot (Morisaki & Raymond, 2020; Veitch et al., 2015b; Pankow et al., 2015; Smith et al., 2020; Berry et al., 2015; Del Pozzo et al., 2018; Romero-Shaw et al., 2020; Biwer et al., 2019; Sidery et al., 2014; Green & Gair, 2021), constructed from $N$ runs on injected signals where the true parameters are known.
@@
A $p$-$p$ plot shows the fraction of events for which the true parameter values lie within a given credible interval as a function of the credible interval. In the ideal case where the parameter inference procedure leads to the correct coverage, this plot should be a diagonal line. There is a statistical uncertainty that arises from the finite number, $N$, of datasets analysed, often indicated by a gray region around the ideal diagonal line (as in Figures 11 and 12), and this is usually the dominant source of uncertainty in $p$-$p$ plot analyses. However, as $N$ increases, the nested sampling uncertainty becomes important. This uncertainty can be evaluated using either the likelihood binning or reconstructed runs method, and we can place a ‘nested sampling confidence interval’ on our $p$-$p$ plot to account for the variation of the calculated percentiles from run to run on the same set of events.
@@
Figures 11 and 12 were produced using the likelihood binning method. For each posterior sample from a given event, the individual parameter values were resampled from bins of neighbouring phantom points. These resampled parameter values were then used to calculate a set of credible intervals for that event, from which the uncertainty on the credible interval was estimated. In Figure 11, a total of $200$ injected signals were analysed, with the injection parameters drawn from the prior, and the gray region shows the 3-$\sigma$ (99.799.7%) confidence interval for $N=200$ due to a finite event sample size. The purple region shows the corresponding 3-$\sigma$ confidence interval on the $p$-$p$ curve due to the nested sampling parameter estimation uncertainty. In scenarios where the $p$-$p$ plot lies outside of the gray region, using the nested sampling error bars may help to assess how much of the deviation is simply due to the stochastic nature of nested sampling. Figure 12 is produced from $1000$ injected signals, where now the statistical uncertainty for $N=1000$ and the nested sampling parameter estimation uncertainty are of similar sizes. Here, it is especially useful to consider the latter uncertainty in assessing to what extent the $p$-$p$ plot displays the expected coverage. Ideally, the gray confidence intervals should include both sources of uncertainties.
@@
Figure 11: The $p$-$p$ plot obtained for the chirp mass using bilby and PolyChord is shown. As with the parameter mean estimates, we only sample over four parameters, simply setting the rest to the injected value. The plot shows the cumulative distribution function (CDF) of the percentiles of the true chirp mass for $200$ events, calculated in the usual way, with the associated 3-$\sigma$ confidence interval due to the statistical uncertainty from the finite number of events (gray). The purple region shows the 3-$\sigma$ confidence interval on the calculated CDF of the percentiles due to the stochastic nature of nested sampling. The binomial error still dominates, but the nested sampling parameter estimation uncertainty can be useful in assessing the source of deviations from the expected distribution.
@@
Figure 12: The CDF of the percentiles of the true chirp mass is shown for $1000$ events, along with the 3-$\sigma$ confidence interval due to a finite event sample size (gray). The 3-$\sigma$ confidence interval due to the nested sampling parameter estimation uncertainty is also shown (purple). These are now of a comparable size, and in order to make an accurate assessment of the coverage, it is necessary to account for both sources of uncertainties.